8(h-1)=6h+4+2h
Solve for h
Please explain the steps, the correct answer is no solution but i want to know how you got that answer.

Answer by Jacky C(3)

8(h-1)=6h+4+2h
since h is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the quation
6h+4+2h=8(h-1)
6h and 2h are like terms, so that means 2h + 6h = 8h
8h+4=8(h-1)
then multiply 8 by each term inside the parentheses.
8h+4=(8h-8)
substracting 8h from both sides
8h+4-8h=-8
-8h and 8h are like terms means -8h+8h=0
0+4=-8
remove the 0
4=-8
since 4 equal to -8 there are no solutions